OpenStudy (anonymous):

A train starts from O and reaches station X with a uniform speed of 160 km/hr and returns back to Owith a uniform speed of 120km/hr. find its average speed and average velocity.

OpenStudy (anonymous):

You can calculate the whole distance traveled by train and then divide it by time it spent to this trip, to get its average speed. Average velocity is zero because the displacement for the whole trip is zero in this case.

OpenStudy (anonymous):

plzz tell the solving

OpenStudy (anonymous):

Distance from O to X = 160 * T1 Distance from X to O = 120 * T2 You can find T1/T2 ratio using the fact that Distance from O to X = Distance from X to O now Average speed = (160 T1+ N T1)/(NT1+T1) Where N is some coefficient is the ratio itself

OpenStudy (anonymous):

Sorry there was mistake Average speed = (160 T1+ 120N T1)/(NT1+T1)